Let T

 be a function from R

3

→R

3

 defined by

T(x,y,z)=(x−y+2z,2x+y,−x−2y+2z)


(i)

(ii)

  Show that T is a Linear Transformation

  Find nullity of T



1
Expert's answer
2022-02-01T12:09:19-0500

A transformation T is linear if


1) T(ax)= aT(x)


2) T(x)=A(x) where A is a matrix.


Part 1


Ta(x,y,z)"=" T(ax,ay,az)

"=(ax-ay+2az,2ax+ay,-ax-2ay+2az)"


"=a(x-y+2z,2x+y,-x-2y+2z)"

"=aT(x,y,z)"



"T(x,y,z)=x\begin{pmatrix}
 1 \\
 2\\
-1
\end{pmatrix}+y\begin{pmatrix}
 -1 \\
 1 \\
-2
\end{pmatrix}+z\begin{pmatrix}
 2 \\
 0 \\
2
\end{pmatrix}"




"T(x,y,z)=\begin{pmatrix}
 1&-1&2 \\
 2&1&0 \\
-1&-2&2
\end{pmatrix}" "\begin{pmatrix}
 x \\
 y\\
z
\end{pmatrix}"


Therefore T is linear


Part 2


"\frac{1}{3}(R_2-2R_1\to\>R_2)"

"R_3+R_1\to\>R_3"



"\begin{pmatrix}
 1&-1& 2\\
 1&1 & \frac{-4}{3}\\
0&-3&4
\end{pmatrix}"


"R_1+R_2\to\>R_1"

"R_3+3R_2\to\>R_3"


"\begin{pmatrix}
 1&0&\frac{2}{3} \\
 0&1&\frac{-4}{3} \\
0&0&0
\end{pmatrix}"


Nullity is "=1"